OpenStudy (anonymous):

find a quadratic equation with roots -1+4i and -1-4i will give medal!!

OpenStudy (anonymous):

you want a real quick method?

OpenStudy (anonymous):

yes that would be nice :)

OpenStudy (anonymous):

if \(a+bi\) are the roots of a quadratic, then the quadratic is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (anonymous):

in your case \[a=1,b=4\] so the quadratic is \[x^2-2\times (-1)x+(1^2+4^2)\] or \[x^2+2x+17\]

OpenStudy (anonymous):

sorry i meant \(a=-1,b=4\) but the answer i wrote is correct if you cannot memorize that, another method is to work backwards \[x=-1+4i\] \[x+1=4i\] \[(x+1)^2=(4i)^2=-16\] \[x^2+2x+1=-16\] \[x^2+2x+17=0\]

OpenStudy (anonymous):

oh! okay thank you :)

OpenStudy (anonymous):

yw easy right?

OpenStudy (anonymous):

i guess i just made it so much harder than it was! :)

OpenStudy (anonymous):

worst method is to try to multiply out \[(x-(-1+4i))(x-(-1-4i))\] you will get stuck somewhere and likely make a mistake